OpenStudy (anonymous):

This term 76 students pass the mathematics test. The number is decreased by 5% of the number of the last term. How many students pass the the test in the last term? This term 76 students pass the mathematics test. The number is decreased by 5% of the number of the last term. How many students pass the the test in the last term? @Mathematics

OpenStudy (asnaseer):

lets call the number of students that passed last terms test 'n'. now, if this term the number that passed fell by 5%, that means this terms total must be 95n/100. so you now solve for:\[\frac{95n}{100}=76\]

OpenStudy (asnaseer):

do you understand?

OpenStudy (anonymous):

how you get 95n?

OpenStudy (asnaseer):

\[95 percent=\frac{95}{100}\]

OpenStudy (asnaseer):

if the number this term was same as last term, then it would be 100% of last terms total and could be written as:\[\frac{100n}{100}=n\]

OpenStudy (asnaseer):

a 5% drop means you are left with (100-5)% = 95%

OpenStudy (anonymous):

Suppose the number of students that passed last year is n. Then 95% of n could be written as: .95n

OpenStudy (anonymous):

Does that part make sense?

OpenStudy (anonymous):

For example, if 100 passed last year, and only 95% of that number passed this year, then .95*100 = 95 passed this year

OpenStudy (anonymous):

Another example, if 20 passed last year, and only 95% of that number passed this year, then .95*20 = 19 passed this year... etc